Change the velocity-time graph to the position-time graph and acceleration-time graph.

10 Fi (5 )
Answer #1

Duning oto » -0.25-(-0.25) one 25) - On/s2 Q = Vprvi 200 - tart ang 2 to 8 = 0.25-(-0.25 » 0.083 m/s2 5-2 bung stolo = 0.25-0